# Efficient estimation of error bounds for quantum multiparametric imaging with constraints

### v2

@@ IV.2 Two-parameter model
$S_{i}(\mathbf{A})=N\eta^{2}G^{(2)}_{ii}=N\eta^{2}\Bigl{|}\sum_{m=1,2}D^{(ii)}_{m}A_{m}^{2}\Bigr{|}^{2},\;i=1,2.$ (34)
For simplicity, we assume that the coefficients $D^{(ii)}_{m}$ are real and symmetric:
$D^{(ii)}_{m}\in\mathcal{R},\;D^{(11)}_{1}=D^{(22)}_{2}=h_{0},\;D^{(11)}_{2}=D^{(22)}_{1}=h_{1}.$ (35)
@@
    where the 2-component likelihood is factorized into two single-component likelihood functions $L(\mathbf{Y}|\mathbf{S}(\mathbf{A}^{\prime}))=L(Y_{1}|S_{1}(\mathbf{A}^{\prime}))L(Y_{2}|S_{2}(\mathbf{A}^{\prime}))$, which are defined by Eq. (29).
@@
Fig. 7 shows the results for the considered 2-parameter model. The estimators listed above are applied to 1000 random samples of the signal $\mathbf{Y}$ for the true values of the transmission amplitudes localized in the lower part (a), the center (b), and the upper part (c) of the physical domain. The estimation results $\mathbf{Y}\mapsto\hat{\mathbf{A}}$ are shown as points in each part of the figure. The statistics of each generated dataset is shown as an ellipse, indicating the mean value $\langle\hat{\mathbf{A}}\rangle$ and the covariance $C=\operatorname{Cov}(\hat{\mathbf{A}},\hat{\mathbf{A}})$ according to the approach described in Appendix F:
@@
Figure 7: Transmission amplitudes estimation for the 2-parameter model. The true parameters values are $A_{1}=A_{2}=0.2$ (a), 0.5 (b), and 0.9 (c), as indicated by black dots. Blue crosses and orange circles show the parameter inference results of a randomly sampled signal for MLE and Bayesian estimate respectively. 1000 samples were used for statistics estimation; 300 samples are shown in the figure; some results coincide due to discretness of the signal. The modeling parameters are $\eta=0.7$, $h_{0}=1$, $h_{1}=0.8$; $N=1000$ (a,b) and 50 (c). Ellipses (blue dotted for MLE and purple dot-dashed for Bayesian estimate) are centered at the mean of the sampled results and have the shape defined by the corresponding covariance matrix (Appendix F). Black dotted ellipse (extending beyond the plotted regions) shows the predicted results scattering according to CRB with the standard FIM $F(\mathbf{A})$. Black solid ellipse corresponds to the regularized and corrected FIM $\tilde{F}(\mathbf{A})$. Straight black lines show the constraints imposed on the parameters $A_{1}$ and $A_{2}$.
The results of random sampling are compared with the error predictions based on CRB and FIM: the standard one $F(\mathbf{A})$, calculated according to Eq. (38), and the FIM $\tilde{F}(\mathbf{A})$ obtained by applying the proposed regularization and correction procedures to $F(\mathbf{A})$. The FIMs $F(\mathbf{A})$ and $\tilde{F}(\mathbf{A})$ are also visualized as ellipses (Appendix F).
@@ IV.3 Practical multiparametric models in quantum imaging
The incident photon pairs are assumed to be ideally correlated, and the signal is described by Eq. (15) with $i=j$ and the sampling points corresponding to the step $d/2$ in the object plane:
$S_{j}(\mathbf{A})=NG^{(2)}(x_{j},x_{j})=N\left|\sum_{m}D^{(jj)}_{m}A_{m}^{2}\right|^{2},$ (43)
@@
Fig. 8(c) shows the dependence of the estimation error $\Delta^{2}$ on the spatial scale of the problem $d/d_{\text{R}}$ for the model object presented in Fig 8(b) with $A_{\text{min}}=0.9$. The number of the detected coincidence events is taken to be $N=10^{4}$. The error $\Delta^{2}$ calculated for the standard FIM $F(\mathbf{A})$ (solid line) accurately describes the variance for unconstrained estimation (red star-marked points). If the constraint $0\leq A_{m}\leq 1$ is imposed during the transmission amplitudes inference according to Eq. (45), the variance (green circles) and MSE (purple crosses) become smaller. Such error decrease is accurately reproduced by using CRB with the corrected FIM $\tilde{F}(\mathbf{A})$ (dashed blue line).
Fig. 8(d) presents similar results for the case $A_{\text{min}}=0$, which requires regularization of FIM. Application of CRB to the standard FIM does not yield meaningful results in that case since $F(\mathbf{A})$ is singular. For that reason, only the constrained estimation was sampled and compared to the error prediction by the regularized and corrected FIM $\tilde{F}(\mathbf{A})$. One can see that the agreement between the results of the proposed approach and direct sampling is good.
@@
$S_{ij}(\mathbf{A})=NG^{(2)}(x_{i},x_{j})=N\left|\sum_{m,l}D^{(ij)}_{ml}A_{m}A_{l}\right|^{2}$ (47)
where the expressions for calculation of the coefficients $D^{(ij)}_{ml}$ are provided in Ref. [28], and $N$ is the expectation value of the total number of detected coincidence events. As previously, we assume independent Poissonian statistics for the components $Y_{ij}$ of the measured signal $\mathbf{Y}$. Therefore, FIM is described by Eq. (8) and takes the form
@@
Fig. 9(b) shows the results of resolution estimation via CRB for $N=10^{5}$. The total reconstruction error $\Delta^{2}$, defined by Eq. (9), is calculated for the standard FIM $F(\mathbf{A})$ (dashed blue line) and the regularized and corrected one $\tilde{F}(\mathbf{A})$ according to the proposed approach (solid orange line). The achievable resolution can be defined as the minimal problem scale (pixel size) $d=d_{\text{min}}$, for which the reconstruction error $\Delta^{2}$ remains within certain reasonable limit: $\Delta^{2}\leq\Delta^{2}_{\text{threshold}}$ [34]. For illustration purposes, the threshold value $\Delta^{2}_{\text{threshold}}=0.1$, consistent with the actual experimental noise in Ref. [28], is used in Fig. 9(b). For such selection of the threshold, the theoretical resolution limit for the standard FIM is $d_{\text{min}}/d_{\text{R}}\approx 0.42$, which is larger than the spatial scale $d/d_{\text{R}}=8.51\text{ $\mu$m}/30.0\text{ $\mu$m}=0.28$ of the problem, successfully solved experimentally [28]. FIM correction according to the proposed approach shifts the resolution limit to the value $d_{\text{min}}/d_{\text{R}}\approx 0.14$, thus, resolving the seeming contradiction of beating the ultimate theoretical limit experimentally.
@@ V Conclusions
The proposed approach was successfully applied to 1-, 2-, 10-, and 24-parameter imaging problems and exhibited a good agreement with other methods of error estimation. For the simplest 1-parameter model, it is possible to describe various estimators and to quantify resolution analytically. Many fundamental theoretical works have already been devoted to such simple models. For 2 parameters, it is still possible to estimate the errors directly, but the calculations are more complicated. When a multiparametric problem is considered, the proposed procedure continues to work well and yields meaningful results, while other approaches require huge computational efforts. In particular, the developed technique of FIM modification helped us to explain the high resolution achieved experimentally in Ref. [28] and to predict the resolution limit for such a constrained problem quantitatively. It is worth noting that the applicability of the developed approach can be extended to problems with a larger number of parameters (e.g., hundreds or thousands) by analyzing sub-problems of a reasonable size after splitting the initial problem into almost independent windows [28].
@@ Appendix A Geometry of likelihood and Fisher information
Fig. 10(a) corresponds to interpretation of FI in terms of the estimation results distribution for a fixed model. If the true value of the parameter equals $\theta$, the possible realizations of the detected signal $Y$ are distributed according to the likelihood $L(Y|\theta)$, shown in the inset (a’) for specific $\theta$ and representing the vertical cross section AB of the map. Inference of the investigated parameter through its MLE $\hat{\theta}$ corresponds to processing the detected signal value $Y$ by mapping $Y\mapsto\hat{\theta}=\operatorname*{argmax}_{\hat{\theta}}L(Y|\hat{\theta})$. Geometrically, each point of the segment AB is mapped onto the curve CD. The resulting distribution $p(\hat{\theta}|\theta)$, shown in the inset (a”), describes frequencies of getting the estimated values $\hat{\theta}$ in a detection-inference procedure if the true value of the investigated parameter equals $\theta$. In the high-SNR limit, the distribution can be approximated by a Gaussian profile (dashed line in inset (a”)). The width of the distribution (total length $l(\text{CA})+l(\text{BD})$ of the segments CA and BD) characterizes the estimation error and is determined by FI.
@@
Appendix B Iterative algorithm of Fisher information matrix modification
@@
Appendix C Examples of width estimation by the second derivative with regularization
@@
Appendix D Fisher information for 1-parameter problem with discrete and continuous signal
@@
First, let us consider the typical case, when the detected signal $Y$ takes integer values distributed according to the Poisson distribution with the mean $S(A)$:
$L_{\text{P}}(Y|S(A))=\frac{S^{Y}(A)}{Y!}e^{-S(A)}.$ (65)
According to Eq. (8), the FI equals

### v3

@@ IV.2 Two-parameter model
$S_{i}(\mathbf{A})=N\eta^{2}G^{(2)}_{ii}=N\eta^{2}\biggl{|}\sum_{m,l=1,2}D^{(ii)}_{ml}A_{m}A_{l}\biggr{|}^{2},\;i=1,2.$ (37)
For simplicity, we assume that the coefficients $D^{(ii)}_{ml}$ are real, symmetric, and correspond to ideal spatial correlations of photons:
$D^{(ii)}_{ml}\in\mathcal{R},\;D^{(11)}_{11}=D^{(22)}_{22}=h_{0},\\ D^{(11)}_{22}=D^{(22)}_{11}=h_{1},\;D^{(ii)}_{12}=D^{(ii)}_{21}=0.$ (38)
@@
    where the likelihood for the 2-component signal is defined according to Eq. (10).
@@
Figure 7 shows the results for the considered 2-parameter model. The estimators listed above are applied to 1000 random samples of the signal $\mathbf{Y}$ for the true values of the transmission amplitudes localized in the lower part (a), the center (b), and the upper part (c) of the physical domain. The estimation results $\mathbf{Y}\mapsto\hat{\mathbf{A}}$ are shown as points in each part of the figure. The statistics of each generated dataset is shown as an ellipse, indicating the mean value $\langle\hat{\mathbf{A}}\rangle$ and the covariance $C=\operatorname{Cov}(\hat{\mathbf{A}},\hat{\mathbf{A}})$ according to the approach described in Appendix H:
@@
Figure 7: Transmission amplitudes estimation for the 2-parameter model. The true parameters values are $A_{1}=A_{2}=0.2$ (a), 0.5 (b), and 0.9 (c), as indicated by black dots. Blue crosses and orange circles show the parameter inference results of a randomly sampled signal for MLE and Bayesian estimate respectively. 1000 samples were used for statistics estimation; 300 samples are shown in the figure; some results coincide due to discretness of the signal. The modeling parameters are $\eta=0.7$, $h_{0}=1$, $h_{1}=0.8$; $N=1000$ (a,b) and 50 (c). Ellipses (blue dotted for MLE and purple dot-dashed for Bayesian estimate) are centered at the mean of the sampled results and have the shape defined by the corresponding covariance matrix (Appendix H). Black dotted ellipse (extending beyond the plotted regions) shows the predicted results scattering according to CRB with the standard FIM $F(\mathbf{A})$. Black solid ellipse corresponds to the regularized and corrected FIM $\tilde{F}(\mathbf{A})$. Straight black lines show the constraints imposed on the parameters $A_{1}$ and $A_{2}$.
The results of random sampling are compared with the error predictions based on CRB and FIM: the standard one $F(\mathbf{A})$, calculated according to Eq. (41), and the FIM $\tilde{F}(\mathbf{A})$ obtained by applying the proposed regularization and correction procedures to $F(\mathbf{A})$. The FIMs $F(\mathbf{A})$ and $\tilde{F}(\mathbf{A})$ are also visualized as ellipses (Appendix H).
@@ IV.3 Practical multiparametric models in quantum imaging
The incident photon pairs are assumed to be ideally correlated, and the signal is described by Eq. (19) with $i=j$, $m=l$, and the sampling points corresponding to the step $d/2$ in the object plane:
$S_{j}(\mathbf{A})=NG^{(2)}(x_{j},x_{j})=N\biggl{|}\sum_{m}D^{(jj)}_{mm}A_{m}^{2}\biggr{|}^{2},$ (46)
@@
Figure 8(c) shows the dependence of the estimation error $\Delta^{2}$ on the spatial scale of the problem $d/d_{\text{R}}$ for the model object presented in Fig 8(b) with $A_{\text{min}}=0.9$. The number of the detected coincidence events is taken to be $N=10^{4}$. The error $\Delta^{2}$ calculated for the standard FIM $F(\mathbf{A})$ (solid line) accurately describes the variance for unconstrained estimation (red star-marked points). If the constraint $0\leq A_{m}\leq 1$ is imposed during the transmission amplitudes inference according to Eq. (48), the variance (green circles) and MSE (purple crosses) become smaller. Such error decrease is accurately reproduced by using CRB with the corrected FIM $\tilde{F}(\mathbf{A})$ (dashed blue line).
Figure 8(d) presents similar results for the case $A_{\text{min}}=0$, which requires regularization of FIM. Application of CRB to the standard FIM does not yield meaningful results in that case since $F(\mathbf{A})$ is singular. For that reason, only the constrained estimation was sampled and compared to the error prediction by the regularized and corrected FIM $\tilde{F}(\mathbf{A})$. One can see that the agreement between the results of the proposed approach and direct sampling is good.
@@
$S_{ij}(\mathbf{A})=NG^{(2)}(x_{i},x_{j})$ (50)
where $G^{(2)}(x_{i},x_{j})$ is calculated according to Eq. (19), the expressions for the coefficients $D^{(ij)}_{ml}$ are provided in Ref. [28], and $N$ is the expectation value of the total number of detected coincidence events. As previously, we assume independent Poissonian statistics for the components $Y_{ij}$ of the measured signal $\mathbf{Y}$. Therefore, FIM is described by Eq. (12) and takes the form
@@
Figure 9(b) shows the results of resolution estimation via CRB for $N=10^{5}$. The total reconstruction error $\Delta^{2}$, defined by Eq. (13), is calculated for the standard FIM $F(\mathbf{A})$ (dashed blue line) and the regularized and corrected one $\tilde{F}(\mathbf{A})$ according to the proposed approach (solid orange line). The achievable resolution can be defined as the minimal problem scale (pixel size) $d=d_{\text{min}}$, for which the reconstruction error $\Delta^{2}$ remains within certain reasonable limit: $\Delta^{2}\leq\Delta^{2}_{\text{threshold}}$ [34]. For illustration purposes, the threshold value $\Delta^{2}_{\text{threshold}}=0.1$, consistent with the actual experimental noise in Ref. [28], is used in Fig. 9(b). For such selection of the threshold, the theoretical resolution limit for the standard FIM is $d_{\text{min}}/d_{\text{R}}\approx 0.42$, which is larger than the spatial scale $d/d_{\text{R}}=8.51\text{ $\mu$m}/30.0\text{ $\mu$m}=0.28$ of the problem, successfully solved experimentally [28]. FIM correction according to the proposed approach shifts the resolution limit to the value $d_{\text{min}}/d_{\text{R}}\approx 0.14$, thus, resolving the seeming contradiction of beating the ultimate theoretical limit experimentally.
@@ V Conclusions
The proposed approach was successfully applied to 1-, 2-, 10-, and 24-parameter imaging problems and exhibited a good agreement with other methods of error estimation. For the simplest 1-parameter model, it is possible to describe various estimators and to quantify resolution analytically. Many fundamental theoretical works have already been devoted to such simple models. For 2 parameters, it is still possible to estimate the errors directly, but the calculations are more complicated. When a multiparametric problem is considered, the proposed procedure continues to work well and yields meaningful results, while other approaches require huge computational efforts. In particular, the developed technique of FIM modification helped us to explain the high resolution achieved experimentally in Ref. [28] and to predict the resolution limit for such a constrained problem quantitatively. While the provided illustrations assumed Poisson distribution of the signal noise and applicability of Eq. (12), the proposed method itself is not limited to such models in any way and can be applied to FIM regardless of the specific type of expression used for its calculation. It is worth noting that the applicability of the developed approach can be extended to problems with a larger number of parameters (e.g., hundreds or thousands) by analyzing sub-problems of a reasonable size after splitting the initial problem into almost independent windows [28].
@@ Appendix A Geometry of likelihood and Fisher information
Figure 10(a) corresponds to interpretation of FI in terms of the estimation results distribution for a fixed model. If the true value of the parameter equals $\theta$, the possible realizations of the detected signal $Y$ are distributed according to the likelihood $L(Y|\theta)$, shown in the inset (a’) for specific $\theta$ and representing the vertical cross section AB of the map. Inference of the investigated parameter through its MLE $\hat{\theta}$ corresponds to processing the detected signal value $Y$ by mapping $Y\mapsto\hat{\theta}=\operatorname*{argmax}_{\hat{\theta}}L(Y|\hat{\theta})$. Geometrically, each point of the segment AB is mapped onto the curve CD. The resulting distribution $p(\hat{\theta}|\theta)$, shown in the inset (a”), describes frequencies of getting the estimated values $\hat{\theta}$ in a detection-inference procedure if the true value of the investigated parameter equals $\theta$. In the high-SNR limit, the distribution can be approximated by a Gaussian profile (dashed line in inset (a”)). The width of the distribution (total length $l(\text{CA})+l(\text{BD})$ of the segments CA and BD) characterizes the estimation error and is determined by FI.
@@
Appendix B Fisher information for independent events
For the specific case of the likelihood, defined by Eq. (10), the general expression (4) can be simplified in the following way. Due to additivity of FI [54] and independence of the signal components (factorization of the expression (10) for the likelihood), the FIM is a sum of separate contributions $F_{\mu\nu}^{(i)}$:
$F_{\mu\nu}=\sum_{i}F_{\mu\nu}^{(i)}$ (53)
where a term
$F_{\mu\nu}^{(i)}=\sum_{Y_{i}=0}^{\infty}\frac{1}{L(Y_{i}|S_{i}(\boldsymbol{\theta}))}\frac{\partial L(Y_{i}|S_{i}(\boldsymbol{\theta}))}{\partial\theta_{\mu}}\frac{\partial L(Y_{i}|S_{i}(\boldsymbol{\theta}))}{\partial\theta_{\nu}}$ (54)
represents the FI for the $i$-th component of the signal.
For the single-component likelihood, described by Eq. (11), the derivatives over the parameters can be expressed as
$\frac{\partial L(Y_{i}|S_{i}(\boldsymbol{\theta}))}{\partial\theta_{\mu}}=\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\mu}}\left(\frac{Y_{i}}{S_{i}(\boldsymbol{\theta})}-1\right)L(Y_{i}|S_{i}(\boldsymbol{\theta})).$ (55)
Substitution of Eq. (55) into Eq. (54) yields the resulting expression
$F_{\mu\nu}^{(i)}=\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\mu}}\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\nu}}\sum_{Y_{i}=0}^{\infty}\left(\frac{Y_{i}}{S_{i}(\boldsymbol{\theta})}-1\right)^{2}L(Y_{i}|S_{i}(\boldsymbol{\theta}))\\ =\frac{1}{S_{i}(\boldsymbol{\theta})}\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\mu}}\frac{\partial S_{i}(\boldsymbol{\theta})}{\partial\theta_{\nu}}.$ (56)
Eq. (12) is obtained by substitution of expression (56) into Eq. (53).
The presented approach to construction of FIM is based on treating the measured dataset as a whole and considering the multivariate probability distribution for the signal vector $\mathbf{Y}$. An alternative approach, typical for quantum imaging [27, 29, 30, 31], is based on calculation of FI $F^{\text{(event)}}$ for a single detection event with subsequent multiplication of the result by the number of detected events $N$. A single detection event is characterized by the probability distribution $p_{i}(\boldsymbol{\theta})$, where the index $i$ enumerates the possible outcomes. For example, when just field intensity is measured ($n=1$ in Eq. (8)), the index $i$ describes the position of a single-photon detection $\mathbf{r}^{(i)}$. The probabilities are normalized as $\sum_{i}p_{i}(\boldsymbol{\theta})=1$, and the FIM can be constructed according to Eq. (4) as
$F_{\mu\nu}=NF_{\mu\nu}^{\text{(event)}}=N\sum_{i}\frac{1}{p_{i}(\boldsymbol{\theta})}\frac{\partial p_{i}(\boldsymbol{\theta})}{\partial\theta_{\mu}}\frac{\partial p_{i}(\boldsymbol{\theta})}{\partial\theta_{\nu}}.$ (57)
One can easily relate the single-event probability distribution $p_{i}(\boldsymbol{\theta})$ with the mean number $S_{i}(\boldsymbol{\theta})$ of such events used in our approach as
$S_{i}(\boldsymbol{\theta})=Np_{i}(\boldsymbol{\theta}).$ (58)
Therefore, Eqs. (12) and (57) yield exactly the same result. Such equivalence of the approaches is not surprising since the Poisson distribution assumed during derivation of Eq. (12) is directly related to the independence of the events used in Eq. (57).
Appendix C Iterative algorithm of Fisher information matrix modification
@@
Appendix D Examples of width estimation by the second derivative with regularization
@@
Appendix E Fisher information for 1-parameter problem with discrete and continuous signal
@@
First, let us consider the typical case, when the detected signal $Y$ takes integer values distributed according to the Poisson distribution (11) with the mean $S(A)$. According to Eq. (12), the FI equals